A bag contains 2 green, 9 brown, 7 yellow and 4 blue marbles. once a marble is selected, it is not replaced. find each probability.
1)p(brown, then yellow)
3)P(yellow, then yellow)
5 P(green, then not blue)

Answer by checkley77(12844)

9/22*7/21=63/462=.136=13.6% chance of a brown then a yellow selection.
7/22*6/21=42/462=.09=9% chance of a yellow then a yellow selection.
2/22*17/21=34/462=.0736=7.36% chance of a green & then not a blue.